state the slope and y-intercept for x+2y=8
Answer by stanbon(75887)

state the slope and y-intercept for x+2y=8
---------------------------
Solve for "y":
2y = -x + 8
y = (-1/2)x + 4
-----
slope = -1/2
y-intercept = 4
====================================
Cheers,
stan H.